OpenStudy (nice):

how to rewrite a triple integral after changing limits ? ( from dxdzdy to dydxdz ) I have the question ...

OpenStudy (nice):

second question in the attached file

OpenStudy (anonymous):

What is all that writing on the problem. Is that your attempt and you want us to check if it is right?

OpenStudy (nice):

no no it's the right answer, but I don't understand how !

OpenStudy (nice):

I need explanation ,, my exam is tomorrow :S :S :S

OpenStudy (amistre64):

1 or 2? or both?

OpenStudy (nice):

2 only ,,

OpenStudy (amistre64):

its all about translating what you got into a new direction to match the switch

OpenStudy (nice):

is there any strategy to get the right answer ?

OpenStudy (amistre64):

well, the outer one is a constant; moves from zero to it high point for starters; z = [0,4] \[\int_{0}^{4}dz\]

OpenStudy (amistre64):

the middle should be in terms of the outer... so it gets a z spot right?

OpenStudy (nice):

I think that I have always to draw ,, I think it's the only way !! but I was hopping to find simpler way ..

OpenStudy (amistre64):

simpler? maybe, but i tend to only now the hard way lol

OpenStudy (nice):

what are you studying ?

OpenStudy (amistre64):

whatever I can get my hands on :)

OpenStudy (nice):

yeah I mean simpler than drawing the graph

OpenStudy (amistre64):

ive taught myself all this stuff; and as i go thru the college courses I learn ways that i was to stupid to pick up on me own

OpenStudy (nice):

aha!! I'm suffering from my doctor in calculus this course, and I'm lost! So NOW I learned that I have to be independent specially on college,, right ?

OpenStudy (nice):

in college *

OpenStudy (amistre64):

dxdzdy dydxdz -------- -------- x: 4-2y-z y: (4-x-z)/2 x: 0 y: 0 z: 4-2y x: 4-z z: 0 x: 0 y: 2 z: 4 y: 0 z: 0 the first is a translation from point to plane x = 4-2y-z <=> y = (4-x-z)/2

OpenStudy (amistre64):

the second is a translation from the seems to keep the inner one ignored x = 4-2y-z (ignore the -2y from the inner) x=4-z

OpenStudy (amistre64):

same with the last? ignore the inners as 0? z = 4 -x translates to z=4

OpenStudy (amistre64):

but thats just a cursory view and i aint got nuthin to prove its a general rule

OpenStudy (amistre64):

maybe if we get some smarter than mes to verify or deny it :)

OpenStudy (amistre64):

see if it works on double integrals maybe....

OpenStudy (nice):

Thanks allot! I will see ...